OpenStudy (anonymous):

Part A: Factor 3x2y2 − 2xy2 − 8y2. Show your work. (4 points) Part B: Factor x2 + 10x + 25. Show your work. (3 points) Part C: Factor x2 − 36. Show your work. (3 points)

OpenStudy (anonymous):

i need steps i have answers! Part A.) y^2(3x+4)(x-2) Part B.) (x-5)^2 Part C.) (x+6)(x-6)

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

sorry I cant help u on this but I love the name :)

OpenStudy (anonymous):

ty :}

OpenStudy (anonymous):

yw and good luck on ur question!

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw

zepdrix (zepdrix):

\[\large\rm 3x^2y^2 − 2xy^2 − 8y^2\]What are we doing? :U Factoring this thing?

zepdrix (zepdrix):

\[\large\rm 3x^2\color{orangered}{y^2} − 2x\color{orangered}{y^2} − 8\color{orangered}{y^2}\]Hmm there isn't a whole lot that we can do with this first one. Notice that the last term has no x's, so we can't pull out any factor out any amount of x. The last two terms have a 2 in them (8 is 4 times 2) but the first term does not, so we can't pull a 2 out. You can do something with the y's though, ya?

OpenStudy (anonymous):

Yes.

zepdrix (zepdrix):

Oh I guess we can factor the x's after that though :3 hah

zepdrix (zepdrix):

\[\large\rm \color{orangered}{y^2}(3x^2−2x−8)\]So you ok with that first step? :o

OpenStudy (anonymous):

Um (y2)(3x2+−2x+−8) =(y2)(3x2)+(y2)(−2x)+(y2)(−8) =3x2y2−2xy2−8y2 =3x2y2−2xy2−8y2

zepdrix (zepdrix):

You just un-did the first step :) Not sure why hehe

OpenStudy (anonymous):

Opps >_<

OpenStudy (anonymous):

I am confused on this..

zepdrix (zepdrix):

So we're taking y^2 out of each term.